6.80 g of soduim chloride are added to 265 g of water. find the mole fraction of thesoduim chloride and water in the solutiom

1
Expert's answer
2022-02-24T10:07:03-0500

Molar mass of NaCl"=58.44g\/mol"

Molar mass of Water "=18.02g\/mol"


Moles of NaCl"=\frac{6.80}{58.44}=0.1164"


Moles of H2O "=\frac{265}{18.02}=14.7059"



Total moles "=14.8223"


Mole fraction of H2O"=\frac{14.7059}{14.8223}=0.9922"


Mole fraction of NaCl "=1-0.9922"

"=0.0078"